Find two level curve of f(x,y)=2xy/(x^2+y^2 ) . Give a rough sketch



1
Expert's answer
2021-06-11T10:10:13-0400
"f(x, y)=\dfrac{2xy}{x^2+y^2}"

"z=\dfrac{2xy}{x^2+y^2}"


"c=0: \dfrac{2xy}{x^2+y^2}=0"


"x=0, y\not=0"

Or

"y=0, x\not=0"




"c=1: \dfrac{2xy}{x^2+y^2}=1"

"x^2-2xy+y^2=0"

"(x-y)^2=0"


"x=y, x\not=0"

Blue graph



"c=-1: \dfrac{2xy}{x^2+y^2}=-1"

"x^2+2xy+y^2=0"

"(x+y)^2=0"


"y=-x, x\not=0"

Violet graph



"c=0.5: \dfrac{2xy}{x^2+y^2}=0.5"

"y^2-4xy+x^2=0"

"y=\dfrac{4x\pm\sqrt{16x^2-4x^2}}{2}""y=(2-\sqrt{3})x\ or\ y=(2+\sqrt{3})x, x\not=0"

Red graph